# Both Stellar Mass and Gravitational Potential Shape the Gas-Phase Metallicity

### v1

@@ Appendix A Properties Included in The Random Forest Regression
$\mathrm{M}_{\mathrm{dyn,R_{e}}}$ dynamical mass from Zhu et al. 2023 within $\mathrm{1R_{e}}$
@@
$\mathrm{\sigma_{e}}$ effective velocity dispersion from Zhu et al. 2023 within $\mathrm{1R_{e}}$
@@ Appendix B Reproducibility of and Comparison to Previous Results
We compare our results with those of Baker & Maiolino (2023) and Sánchez-Menguiano et al. (2024a), both of which applied random forest regression to MaNGA data but reported opposite conclusions. By reproducing their sample selections and parameter sets, we can replicate their main results, allowing a direct comparison with our findings.
@@
Baker & Maiolino (2023) analysed a sample of roughly 1,000 galaxies, less than a quarter of the size of our dataset, and considered only spectroscopic stellar mass measurements from the pyPipe3D catalogue, without including photometric estimates. They concluded that gas-phase metallicity fundamentally depends on the spectroscopic stellar mass. In contrast, Sánchez-Menguiano et al. (2024a) used parameters exclusively from pyPipe3D measured at one effective radius, defining metallicity from a linear fit to the radial profile between $\rm 0.5,R_{\mathrm{e}}$ and $\rm 1.5,R_{\mathrm{e}}$, evaluated at $\rm 1.0,R_{\mathrm{e}}$. Their analysis therefore, omits the central $\rm 0.5,R_{\mathrm{e}}$ region, and they reported the baryonic gravitational potential derived from photometric stellar mass as the primary metallicity driver.
@@
Our results agree with Sánchez-Menguiano et al. (2024a) that photometric stellar mass carries substantially more predictive power for gas-phase metallicity than its spectroscopic counterpart. This likely reflects the higher S/N of broadband photometric SED fits compared to spectroscopy, as well as the tendency of tree-based algorithms like random forest to favour lower-noise, higher-variance features (Kauffmann et al., 2003; Drory et al., 2004; Sánchez-Menguiano et al., 2024a). However, we differ from Sánchez-Menguiano et al. (2024a) in the relative importance of mass versus gravitational potential.
@@
By examining feature importance as a function of galactocentric radius, we find that stellar mass dominates within the inner regions ($\rm\lesssim 1.25,R_{\mathrm{e}}$), while the baryonic gravitational potential becomes more important at larger radii ($\rm>1.25,R_{\mathrm{e}}$). This provides a natural explanation for the differences: because Sánchez-Menguiano et al. (2024a) excluded the central $\rm 0.5\mathrm{R_{e}}$ region (where we find stellar mass to be most influential), their analysis emphasised the gravitational potential. Our approach, which includes the galaxy centre, recovers a stronger mass dependence.
@@
These comparisons highlight that discrepancies between studies arise not only from differences in sample size and parameter choice, but also from the spatial regions over which metallicity is defined. As previous work has shown, metallicity correlations and gradients are highly sensitive to the spatial scale of the measurement (Belfiore et al., 2017, 2019b; Maiolino & Mannucci, 2019; Sharda et al., 2024; Garcia et al., 2025).
Appendix C The Radial Dependence in Detail
In this section, we explore the radial dependence discussed in section 4.1.2 where we pointed out that there exists a radial dependence of the gas-phase metallicity: at higher radii ($\rm>1.25R_{e}$) the metallicity is dominated by $\mathrm{\Phi_{*,phot}}$, while at lower radii ($\mathrm{<1.25R_{e}}$) $\mathrm{M_{*,phot}}$ is the key driver. Here, we repeat the same analysis, but instead of computing the metallicity within widening radii, we calculate it within annuli at different radii. Figure 8 shows our results, which are almost identical to the version presented in Figure 2, but with an even stronger dependence on $\mathrm{\Phi_{*,phot}}$ within the high-radii annuli.
In this section, we revisit the radial dependence discussed in Section 4.1.2, where we showed that gas-phase metallicity varies with radius: at larger radii ($\rm>1.25R_{e}$), metallicity is primarily governed by $\mathrm{\Phi_{*,phot}}$, whereas at smaller radii ($\rm<1.25R_{e}$), $\mathrm{M_{*,phot}}$ is the dominant driver. Here, instead of computing metallicity within cumulative apertures of increasing radius, we calculate it within annuli at different radial intervals. Figure 8 presents the results, which closely resemble those in Figure 2, but with an even stronger dependence on $\mathrm{\Phi_{*,phot}}$ in the outer annuli.
Figure 8: Same as Figure 2, except here the metallicity is averaged within concentric annuli at fixed radii, as indicated on the x-axis, rather than within cumulative apertures of increasing radius.
Appendix D Differential Measurement Uncertainty Test
@@
To look into this in detail, we additionally perform a differential measurement uncertainty test by artificially adding Gaussian random noise in units of the uncertainty of $\mathrm{M_{dyn,R_{e}}}$ ($0.24$ dex, see Campbell et al., 2017) to $\mathrm{M_{*,phot}}$ and re-evaluate their importances in driving the gas-phase metallicity via a random forest regression. Figure 9 shows the relative importances of $\mathrm{M_{*,phot}}$, uniform random variable, and $\mathrm{M_{dyn,R_{e}}}$ in driving the gas-phase metallicity for the cases of no noise and different noise levels being added to the spectro-photometric stellar mass. We find that by adding $\mathrm{\geq 3\sigma_{M_{dyn,R_{e}}}}$ Gaussian random noise to $\mathrm{M_{*,phot}}$ the importance of $\mathrm{M_{*,phot}}$ starts to diminish and $\mathrm{M_{dyn,R_{e}}}$ becomes the main driver of the metallicity. However, the importance of $\mathrm{M_{dyn,R_{e}}}$ doesn’t fully reach the high importance of $\mathrm{M_{*,phot}}$ even with an additional noise of $\mathrm{9\sigma_{M_{dyn,R_{e}}}}$. Additionally, adding $\mathrm{9\sigma_{M_{dyn,R_{e}}}}$ levels of noise results in $\mathrm{M_{*,phot}}$ being as (un-)important as a random variable.
@@
Appendix E Combined Effect of Stellar Mass and Effective Radius
We include an additional random forest analysis by conducting a similar investigation into the dependence of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ as Sánchez-Menguiano et al. (2024a) has done. Essentially, we investigate the relation between the gas-phase metallicity and $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ for different $\mathrm{\alpha\in[0.1,0.2,...,2.0]}$. For $\mathrm{\alpha=0}$ we establish the stellar mass, for $\mathrm{\alpha=1.0}$ we have the gravitational potential $\mathrm{\Phi_{*}}$, and for $\mathrm{\alpha=2}$ we get the mass surface density $\mathrm{\Sigma_{*}}$. Sánchez-Menguiano et al. (2024a) reasoning behind this exploration is that the correlation between the gas-phase metallicity and both stellar mass and galaxy size has been previously found (Ellison et al., 2008; Sánchez Almeida & Dalla Vecchia, 2018; Sánchez et al., 2017). Within our analysis above, we include $\mathrm{\Phi_{*}=M_{*}/R_{e}}$, whose contribution to the metallicity could be attributed to a combined effect from stellar mass and galaxy size.
We present our results for the MaNGA survey in Figure 10 (top figure), where the magenta data points indicate the importance of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ and the blue points are the stellar mass. For an $\mathrm{\alpha=0}$, we only include the stellar mass once in the random forest; thus, there is no orange datapoint for this. Besides the already confirmed results of the stellar mass being the main driver of the gas-phase metallicity, we also find that there is a significant importance of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ for $\mathrm{\alpha}$ of roughly 0.1 to 0.8, when it dominates over the stellar mass. This is comparable to Sánchez-Menguiano et al. (2024a) results of $\mathrm{\alpha=0.6}$. For $\mathrm{\alpha>0.8}$, we confirm that $\mathrm{M_{*}}$ remains the dominant driver of the gas-phase metallicities.
@@ Appendix E Combined Effect of Stellar Mass and Effective Radius
Appendix F Partial Correlation Coefficients Analysis for SDSS
@@
Appendix G Central versus Satellite Galaxies with SDSS
We use central satellite classifications from Yang et al. (2005, 2007) to identify satellite galaxies among our sample. Figure 13 shows the random forest analysis as done for Figure 3 while separating central and satellite galaxies. We do not find any difference in the results between central and satellite galaxies, as has previously been found for the stellar metallicity by Baker et al. (2024).

### v2

@@ Appendix A Properties Included in The Random Forest Regression
$\mathrm{M}_{\mathrm{dyn,R_{e}}}$ dynamical mass from zhu_manga_2023 within $\mathrm{1R_{e}}$
@@
$\mathrm{\sigma_{e}}$ effective velocity dispersion from zhu_manga_2023 within $\mathrm{1R_{e}}$
@@ Appendix B Reproducibility of and Comparison to Previous Results
We compare our results with those of baker_stellar_2023 and sanchez-menguiano_stellar_2024, both of which applied random forest regression to MaNGA data but reported opposite conclusions. By reproducing their sample selections and parameter sets, we can replicate their main results, allowing a direct comparison with our findings.
@@
baker_stellar_2023 analysed a sample of roughly 1,000 galaxies, less than a quarter of the size of our dataset, and considered only spectroscopic stellar mass measurements from the pyPipe3D catalogue, without including photometric estimates. They concluded that gas-phase metallicity fundamentally depends on the spectroscopic stellar mass. In contrast, sanchez-menguiano_stellar_2024 used parameters exclusively from pyPipe3D measured at one effective radius, defining metallicity from a linear fit to the radial profile between $\rm 0.5,R_{\mathrm{e}}$ and $\rm 1.5,R_{\mathrm{e}}$, evaluated at $\rm 1.0,R_{\mathrm{e}}$. Their analysis therefore, omits the central $\rm 0.5,R_{\mathrm{e}}$ region, and they reported the baryonic gravitational potential derived from photometric stellar mass as the primary metallicity driver.
@@
Our results agree with sanchez-menguiano_stellar_2024 that photometric stellar mass carries substantially more predictive power for gas-phase metallicity than its spectroscopic counterpart. This likely reflects the higher S/N of broadband photometric SED fits compared to spectroscopy, as well as the tendency of tree-based algorithms like random forest to favour lower-noise, higher-variance features (kauffmann_host_2003; drory_comparing_2004; sanchez-menguiano_stellar_2024). However, we differ from sanchez-menguiano_stellar_2024 in the relative importance of mass versus gravitational potential.
@@
By examining feature importance as a function of galactocentric radius, we find that stellar mass dominates within the inner regions ($\rm\lesssim 0.7,R_{\mathrm{e}}$), while the baryonic gravitational potential becomes more important at larger radii ($\rm>0.9,R_{\mathrm{e}}$). This provides a natural explanation for the differences: because sanchez-menguiano_stellar_2024 excluded the central $\rm 0.5\mathrm{R_{e}}$ region (where we find stellar mass to be most influential), their analysis emphasised the gravitational potential. Our approach, which includes the galaxy centre, recovers a stronger mass dependence.
@@
These comparisons highlight that discrepancies between studies arise not only from differences in sample size and parameter choice, but also from the spatial regions over which metallicity is defined. As previous work has shown, metallicity correlations and gradients are highly sensitive to the spatial scale of the measurement (belfiore_sdss_2017; belfiore_bathtub_2019; maiolino_re_2019; sharda_interplay_2024; garcia_metallicity_2025).
Appendix C Differential Measurement Uncertainty Test
@@
To look into this in detail, we additionally perform a differential measurement uncertainty test by artificially adding Gaussian random noise in units of the uncertainty of $\mathrm{M_{dyn,R_{e}}}$ ($0.24$ dex, see campbell_knowing_2017) to $\mathrm{M_{*,phot}}$ and re-evaluate their importances in driving the gas-phase metallicity via a random forest regression. Figure 8 shows the relative importances of $\mathrm{M_{*,phot}}$, uniform random variable, and $\mathrm{M_{dyn,R_{e}}}$ in driving the gas-phase metallicity for the cases of no noise and different noise levels being added to the spectro-photometric stellar mass. We find that by adding $\mathrm{\geq 3\sigma_{M_{dyn,R_{e}}}}$ Gaussian random noise to $\mathrm{M_{*,phot}}$ the importance of $\mathrm{M_{*,phot}}$ starts to diminish and $\mathrm{M_{dyn,R_{e}}}$ becomes the main driver of the metallicity. However, the importance of $\mathrm{M_{dyn,R_{e}}}$ doesn’t fully reach the high importance of $\mathrm{M_{*,phot}}$ even with an additional noise of $\mathrm{9\sigma_{M_{dyn,R_{e}}}}$. Additionally, adding $\mathrm{9\sigma_{M_{dyn,R_{e}}}}$ levels of noise results in $\mathrm{M_{*,phot}}$ being as (un-)important as a random variable.
@@
Appendix D Combined Effect of Stellar Mass and Effective Radius
We include an additional random forest analysis by conducting a similar investigation into the dependence of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ as sanchez-menguiano_stellar_2024 has done. Essentially, we investigate the relation between the gas-phase metallicity and $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ for different $\mathrm{\alpha\in[0.1,0.2,...,2.0]}$. For $\mathrm{\alpha=0}$ we establish the stellar mass, for $\mathrm{\alpha=1.0}$ we have the gravitational potential $\mathrm{\Phi_{*}}$, and for $\mathrm{\alpha=2}$ we get the mass surface density $\mathrm{\Sigma_{*}}$. sanchez-menguiano_stellar_2024 reasoning behind this exploration is that the correlation between the gas-phase metallicity and both stellar mass and galaxy size has been previously found (ellison_clues_2008; sanchez_almeida_origin_2018; sanchez_mass-metallicity_2017). Similarly, ma_revisiting_2024 recently demonstrated that combining stellar mass and size, through $\mathrm{M_{*}/R_{e}^{\alpha}}$ with $\alpha\sim 0.6-1$, provides a much tighter correlation with gas-phase metallicity than stellar mass alone, in both MaNGA observations and TNG50 simulations. Within our analysis above, we include $\mathrm{\Phi_{*}=M_{*}/R_{e}}$, whose contribution to the metallicity could be attributed to a combined effect from stellar mass and galaxy size.
We present our results for the MaNGA survey in Figure 9 (top figure), where the magenta data points indicate the importance of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ and the blue points are the stellar mass. For an $\mathrm{\alpha=0}$, we only include the stellar mass once in the random forest; thus, there is no blue datapoint for this. Besides the already confirmed results of the stellar mass being the main driver of the gas-phase metallicity, we also find that there is a significant importance of $\mathrm{M_{*,phot}/R_{e}^{\alpha}}$ for $\mathrm{\alpha}$ of roughly 0.1 to 0.8, when it dominates over the stellar mass. This is comparable to sanchez-menguiano_stellar_2024 results of $\mathrm{\alpha=0.6}$. For $\mathrm{\alpha>0.8}$, we confirm that $\mathrm{M_{*}}$ remains the dominant driver of the gas-phase metallicities.
@@ Appendix D Combined Effect of Stellar Mass and Effective Radius
Appendix E Partial Correlation Coefficients Analysis for SDSS
@@
Appendix F Central versus Satellite Galaxies with SDSS
We use central satellite classifications from yang_halo-based_2005; yang_galaxy_2007 to identify satellite galaxies among our sample. Figure 12 shows the random forest analysis as done for Figure 3 while separating central and satellite galaxies. We do not find any difference in the results between central and satellite galaxies, as has previously been found for the stellar metallicity by baker_different_2024.